Hyphenation Exception Log

Barbara Beeton

This is the periodic update^{1} of the list of words that TEX fails to hyphenate properly. The list last appeared in full in TUGboat 16, no. 1, starting on page 12, with an update in TUGboat 20, no. 1, pages 50–51. The full list is included here. Owing to the length of the list, it has been subdivided into two parts: English words, and names and non-English words that occur in English texts.

This list is specific to the hyphenation patterns that appear in the original hyphen.tex, that is, the patterns for U.S. English. If such information for other patterns becomes available, consideration will be given to listing that too. (See below, “Hyphenation for languages other than English”.)

In the list below, the first column gives results from TEX’s \showhyphens{...}; entries in the second column are suitable for inclusion in a

\hyphenation{...} list.

In most instances, inflected forms are not shown for nouns and verbs; note that all forms must be specified in a \hyphenation{...} list if they occur in your document.

Thanks to all who have submitted entries to the list. Since some suggestions demonstrated a lack of familiarity with the rules of the hyphenation algorithm, here is a short reminder of the relevant idiosyncrasies. Hyphens will not be inserted before the number of letters specified by \lefthyphenmin, nor after the number of letters specified by \righthyphenmin. For U.S. English,

\lefthyphenmin=2 and \righthyphenmin=3; thus

no word shorter than five letters will be hyphenated. (For the details, see The TEXbook, page 454. For a digression on other views of hyphenation rules, see below under “English hyphenation”.) This particular rule is violated in some of the words listed; however, if a word is hyphenated correctly by TEX except for “missing” hyphens at the beginning or end, it has not been included here.

Some other permissible hyphens have been omitted for reasons of style or clarity. While this is at least partly a matter of personal taste, an author should think of the reader when deciding whether or not to permit just one more break-point in some obscure or confusing word. There really are times when a bit of rewriting is preferable.

One other warning: Some words can be more than one part of speech, depending on context, and have different hyphenations; for example, ‘analyses’ can be either a verb or a plural noun. If such a word appears in this list, hyphens are shown only for the portions of the word that would be hyphenated the same regardless of usage. These words are marked with a ‘*’; additional hyphenation points, if needed in your document, should be inserted with discretionary hyphens.

Words added since the last appearance of the list are preceded by ^{+}.

The reference used to check these hyphenations is Webster’s Third New International Dictionary, Unabridged.

English hyphenation

It has been pointed out to me that the hyphenation rules of British English are based on the etymology of the words being hyphenated as opposed to the “syllabic” principles used in the U.S. Furthermore, in the U.K., it is considered bad style to hyphenate a word after only two letters. In order to make TEX defer hyphenation until after three initial letters,

                                              set \lefthyphenmin=3.

Of course, British hyphenation patterns should be used as well. A set of patterns for UK English has been created by Dominik Wujastyk and Graham Toal, using Frank Liang’s PATGEN and based on a file of 114925 British-hyphenated words generously made available to Dominik by Oxford University Press. (This list of words and the hyphenation break points in the words are copyright to the OUP and may not be redistributed.) The file of hyphenation patterns may be freely distributed; it is posted on CTAN in the file

                                              tex-archive/language/hyphenation/ukhyph.tex

and can be retrieved by anonymous FTP or using a Web browser.

Hyphenation for languages

other than English

Patterns now exist for many languages other than English, including languages using accented alphabets. CTAN holds an extensive collection of

                                              patterns in tex-archive/language/hyphenation

and its subdirectories.

The List — English words

academy(ies)  acad-e-my(ies)
+ ac-cusative  ac-cu-sa-tive
addable  add-a-ble

                                                          ad-di-ble  add-i-ble

adrenaline  adren-a-line
aerospace  aero-space

1 Appeared in TUGboat 21(2), June, 2000 on page

af-terthought  af-ter-thought
agronomist  agron-o-mist
am-phetamine  am-phet-a-mine
anal-yse(d)  an-a-lyse(d)

anal-y-ses  analy-ses *

anomaly(ies)  anom-aly(ies)
an-tideriva-tive  an-ti-deriv-a-tive
anti-nomy(ies)  an-tin-o-my(ies)
antin-u-clear  an-ti-nu-clear
antin-u-cleon  an-ti-nu-cle-on
an-tirev-o-lu-tion-ary  an-ti-rev-o-lu-tion-ary
apotheoses  apoth-e-o-ses
apotheo-sis  apoth-e-o-sis
ap-pendix  ap-pen-dix
archipelago  ar-chi-pel-ago
archive  ar-chive
archiv-ing  ar-chiv-ing
archivist  ar-chiv-ist
archety-pal  ar-che-typ-al
archetyp-i-cal  ar-che-typ-i-cal

arc-t-an-gent  arc-tan-gent

(better: arc tangent)

assignable  as-sign-a-ble

as-sig-nor  as-sign-or

as-sis-tantship  as-sist-ant-ship
asymp-tomatic  asymp-to-matic
asymp-totic  as-ymp-tot-ic
asyn-chronous  asyn-chro-nous
atheroscle-ro-sis  ath-er-o-scle-ro-sis

at-mo-sphere  at-mos-phere

at-tributed  at-trib-uted
at-tributable  at-trib-ut-able
au-toma-tion  au-to-ma-tion
au-toma-ton(a)  au-tom-a-ton(a)
avoirdupois  av-oir-du-pois

ban-dleader  band-leader

bankrupt(cy)  bank-rupt(-cy)

ba-ronies  bar-onies
base-li-neskip             \base-line-skip

bathymetry  ba-thym-e-try
bathyscaphe  bathy-scaphe

bea-nies  bean-ies

be-haviour  be-hav-iour

be-vies  bevies
bib-li-og-ra-phystyle      \bib-li-og-ra-phy-style
bid-if-fer-en-tial  bi-dif-fer-en-tial

biggest  big-gest

bil-l-able  bill-able

biomath-e-mat-ics  bio-math-e-mat-ics
biomedicine  bio-med-i-cine
biorhythms  bio-rhythms
bitmap  bit-map

blan-der  bland-er
blan-d-est  bland-est
blin-der  blind-er
blon-des  blondes

blueprint  blue-print
bornolog-i-cal  bor-no-log-i-cal
bo-tulism  bot-u-lism

brus-quer  brusquer
bus-ier  busier
bus-i-est  busiest
buss-ing  bussing
but-ted  butted

                                                          buz-zword  buzz-word

ca-cophony(ies)  ca-coph-o-ny(ies)

                                                          cam-er-a-men  cam-era-men

cartwheel  cart-wheel

                                                          catar-rhs  ca-tarrhs

catas-trophic  cat-a-stroph-ic
catas-troph-i-cally  cat-a-stroph-i-cally
cauliflower  cau-li-flow-er
cha-parral  chap-ar-ral
chartreuse  char-treuse
cholesteric  cho-les-teric
cigarette  cig-a-rette

                                                          cin-que-foil  cinque-foil

cognac  co-gnac

                                                          cog-nacs  co-gnacs

columns  col-umns
com-parand  com-par-and

                                                          com-para-nds  com-par-ands

com-pendium  com-pen-dium
comptroller  comp-trol-ler
con-formable  con-form-able
con-formist  con-form-ist

                                                          con-for-mity  con-form-ity

congress  con-gress
con-tribute(s,d)  con-trib-ute(s,d)
cose-cant  co-se-cant
cotan-gent  co-tan-gent
courses  cour-ses
crankshaft  crank-shaft
crocodile  croc-o-dile
crosshatch(ed)  cross-hatch(ed)
+ cryp-togram(s)  cryp-to-gram(s)
+ cuneiform  cu-nei-form
dachshund  dachs-hund
database  data-base

                                                          dat-a-p-ath  data-path
                                                          dat-es-tamp  date-stamp

declarable  de-clar-able
defini-tive  de-fin-i-tive
delectable  de-lec-ta-ble
+ demisemi-qua-ver  demi-semi-qua-ver
democratism  de-moc-ra-tism

                                                          de-mos  demos

deriva-tive  de-riv-a-tive
diffract  dif-fract

                                                          di-rer  direr
                                                          di-re-ness  dire-ness

dis-parand  dis-par-and

                                                          dis-traugh-tly  dis-traught-ly

dis-tribute(d)  dis-trib-ute(d)
dou-blespace(ing)  dou-ble-space(-ing)

                                                          dol-lish  doll-ish
                                                          drif-tage  drift-age

drivers  driv-ers
dromedary(ies)  drom-e-dary(ies)
duopolist  du-op-o-list
duopoly  du-op-oly
+ dyslexia  dys-lexia

                                                          eas-t-en-ders  east-end-ers

eco-nomics  eco-nom-ics
economist  econ-o-mist

                                                          elec-trome-chan-i-cal  electro-mechan-i-cal
                                                          elec-tromechanoa-cous-tic electro-mechano-acoustic
                                                          eli-tist  elit-ist

en-trepreneur(ial)  en-tre-pre-neur(-ial)
epinephrine  ep-i-neph-rine

equiv-ari-ant(ce)  equi-vari-ant(ce)

ethy-lene  eth-yl-ene
ev-ersible  ever-si-ble
ev-ert(s,ed,ing)  evert(s,-ed,-ing)
exquisite  ex-quis-ite
ex-traor-di-nary  ex-tra-or-di-nary
fermions  fermi-ons

flag-el-lum(la)  fla-gel-lum(-la)

flammables  flam-ma-bles

fledgling  fledg-ling
flowchart  flow-chart

formidable(y)  for-mi-da-ble(y)
forsythia  for-syth-ia
forthright  forth-right
freeloader  free-loader
friendlier  friend-lier
frivolity  fri-vol-ity
frivolous  friv-o-lous
galac-tic  ga-lac-tic
galaxy(ies)  gal-axy(-ies)
ga-some-ter  gas-om-e-ter
geodesic  ge-o-des-ic
geode-tic  ge-o-det-ic

ge-o-met-ric  geo-met-ric

geotropism  ge-ot-ro-pism
gnomon  gno-mon
grievance  griev-ance
grievous(ly)  griev-ous(-ly)
hairstyle  hair-style
hairstylist  hair-styl-ist

half-s-pace  half-space

harbinger  har-bin-ger
harlequin  har-le-quin
hatcheries  hatch-eries
+ hemidemisemi-qua-ver  hemi-demi-semi-qua-ver
hemoglobin  he-mo-glo-bin
hemophilia  he-mo-phil-ia
hemorhe-ol-ogy  hemo-rhe-ol-ogy
hep-atic  he-pat-ic
hermaphrodite(ic)  her-maph-ro-dite(-ic)
heroes  he-roes

hex-adec-i-mal  hexa-dec-i-mal

holon-omy  ho-lo-no-my

ho-mo-th-etic  ho-mo-thetic

horseradish  horse-rad-ish
hy-potha-la-mus  hy-po-thal-a-mus

ide-als  ideals

ideographs  ideo-graphs

id-iosyn-crasy  idio-syn-crasy

ig-niter  ig-nit-er

ig-n-i-tor  ig-ni-tor
ig-nores-paces             \ignore-spaces

impedances  im-ped-ances
in-finitely  in-fin-ite-ly
in-finites-i-mal  in-fin-i-tes-i-mal

in-fras-truc-ture  in-fra-struc-ture

in-ter-dis-ci-plinary  in-ter-dis-ci-pli-nary
in-ter-galac-tic  in-ter-ga-lac-tic

inu-tile  in-utile

inu-til-ity  in-util-i-ty
+ ir-re-ducible  ir-re-duc-ible

ir-re-vo-ca-ble  ir-rev-o-ca-ble

itinerary(ies)  itin-er-ary(-ies)
jeremi-ads  je-re-mi-ads
keystroke  key-stroke

kil-ning  kiln-ing

la-ciest  lac-i-est
lamentable  lam-en-ta-ble

land-sca-per  land-scap-er

larceny(ist)  lar-ce-ny(-ist)
let-terspac-ing  let-ter-spac-ing

lifes-pan  life-span

lifestyle  life-style
lightweight  light-weight
limousines  lim-ou-sines
linebacker  line-backer

lines-pac-ing              \line-spacing

li-oness  li-on-ess
lithographed  lith-o-graphed
lithographs  lith-o-graphs
lobotomy(ize)  lo-bot-omy(-ize)

lo-ges  loges

longest  long-est
lo-quacity  lo-quac-ity

macroe-co-nomics  macro-eco-nomics

malapropism  mal-a-prop-ism
manuscript  man-u-script
marginal  mar-gin-al
math-e-mati-cian  math-e-ma-ti-cian

mat-tes  mattes
med-i-caid  med-ic-aid

mediocre  medi-ocre
medi-o-crities  medi-oc-ri-ties

me-galith  mega-lith

metabolic  meta-bol-ic
metabolism  me-tab-o-lism

met-a-lan-guage  meta-lan-guage

metaphoric  meta-phor-ic
metropo-lis(es)  me-trop-o-lis(es)
metropoli-tan  met-ro-pol-i-tan

mi-croe-co-nomics  micro-eco-nomics
mi-crofiche  mi-cro-fiche
mil-lage  mill-age

milliliter  mil-li-liter
mimeographed  mimeo-graphed
mimeographs  mimeo-graphs
mimi-cries  mim-ic-ries

mi-nis  min-is
min-isym-po-sium(a)  mini-sym-po-sium(a)

min-uter(est)  mi-nut-er(-est)
mis-chievously  mis-chie-vous-ly

mis-ers  mi-sers

mis-ogamy  mi-sog-a-my
mod-elling  mod-el-ling
molecule  mol-e-cule

monar-chs  mon-archs
mon-eylen-der  money-len-der

monochrome  mono-chrome

mo-noen-er-getic  mono-en-er-getic

monoid  mon-oid
monopole  mono-pole
monopoly  mo-nop-oly

monos-pline  mono-spline
monos-trofic  mono-strofic

mono-tonies  mo-not-o-nies
monotonous  mo-not-o-nous

mo-ro-nism  mo-ron-ism

mosquito  mos-qui-to

mu-d-room  mud-room

mul-ti-faceted  mul-ti-fac-eted

mul-ti-pli-ca-ble  mul-ti-plic-able

mul-tiuser  multi-user (better

with explicit hyphen)

ne-ofields  neo-fields
neon-azi  neo-nazi

newslet-ter  news-let-ter

non-ame  no-name
none-mer-gency  non-emer-gency

nonequiv-ari-ance  non-equi-vari-ance
nonethe-less  none-the-less
noneu-clidean  non-euclid-ean

non-i-so-mor-phic  non-iso-mor-phic
nonpseu-do-com-pact  non-pseudo-com-pact
non-s-mooth  non-smooth

nonuni-form(ly)  non-uni-form(-ly)
nore-pinephrine  nor-ep-i-neph-rine
notwith-stand-ing  not-with-stand-ing
nutcracker  nut-crack-er

oer-st-eds  oer-steds

oligopolist  oli-gop-o-list
oligopoly(ies)  oli-gop-oly(-ies)
operand(s)  op-er-and(s)
orangutan  orang-utan
or-thodon-tist  or-tho-don-tist

or-thok-er-a-tol-ogy  or-tho-ker-a-tol-ogy
or-thoni-tro-toluene  ortho-nitro-toluene

(or-tho-ni-tro-tol-u-ene)

overview  over-view

ox-i-dic  ox-id-ic

padding  pad-ding
painlessly  pain-less-ly

pal-mate  palmate

parabola  par-a-bola
parabolic  par-a-bol-ic
paraboloid  pa-rab-o-loid
paradigm  par-a-digm
parachute  para-chute
paradimethyl-ben-zene  para-di-methyl-benzene

(para-di-meth-yl-ben-zene)

paraflu-o-ro-toluene  para-fluoro-toluene

(para-flu-o-ro-tol-u-ene)

para-g-ra-pher  para-graph-er
par-ale-gal  para-le-gal

par-al-lelism  par-al-lel-ism

para-m-ag-netism  para-mag-net-ism

paramedic  para-medic

param-ethy-lanisole  para-methyl-anisole

(para-meth-yl-an-is-ole)

parametrize  pa-ram-e-trize
paramil-i-tary  para-mil-i-tary
paramount  para-mount
pathogenic  path-o-gen-ic
pee-vish(ness)  peev-ish(-ness)
pen-tagon  pen-ta-gon
petroleum  pe-tro-le-um
phe-nomenon  phe-nom-e-non
philatelist  phi-lat-e-list
phoneme  pho-neme
phone-mic  pho-ne-mic

phos-pho-ric  phos-phor-ic

pho-tographs  pho-to-graphs
pi-cador  pic-a-dor

pi-ran-has  pi-ra-nhas
pla-ca-ble  placa-ble
plea-sance  pleas-ance

poltergeist  pol-ter-geist
polyene  poly-ene
polyethy-lene  poly-eth-yl-ene
polygamist(s)  po-lyg-a-mist(s)
poly-go-niza-tion  polyg-on-i-za-tion
polyphonous  po-lyph-o-nous
polystyrene  poly-styrene
pomegranate  pome-gran-ate

poroe-las-tic  poro-elas-tic

porous  por-ous
postam-ble  post-am-ble
postscript  post-script

pos-tu-ral  pos-tur-al

pream-ble  pre-am-ble
preloaded  pre-loaded
prepar-ing  pre-par-ing
preprint(s)  pre-print(s)

pre-pro-ces-sor  pre-proces-sor
pre-s-plit-ting            \pre-split-ting

priestesses  priest-esses
pret-typrinter(ing)  pret-ty-prin-ter(ing)

pro-ce-du-ral  pro-ce-dur-al
pro-cess  process*

procu-rance  pro-cur-ance

pro-ge-nies  prog-e-nies

progeny  prog-e-ny
pro-hibitive(ly)  pro-hib-i-tive(-ly)
prosci-utto  pro-sciut-to
protester(s)  pro-test-er(s)
protestor(s)  pro-tes-tor(s)
pro-tolan-guage  pro-to-lan-guage

pro-to-ty-pal  pro-to-typ-al

+ province(s)  prov-ince(s)
+ provin-cial  pro-vin-cial

pseu-dod-if-fer-en-tial  pseu-do-dif-fer-en-tial
pseud-ofi-nite  pseu-do-fi-nite

pseud-ofinitely  pseu-do-fi-nite-ly

pseud-o-forces  pseu-do-forces

pseudonym  pseu-do-nym
pseu-doword  pseu-do-word
psychedelic  psy-che-del-ic

psy-chs  psychs

pubescence  pu-bes-cence
quadding  quad-ding
quadratic(s)  qua-drat-ic(s)
quadra-ture  quad-ra-ture
quadriplegic  quad-ri-pleg-ic
quain-ter(est)  quaint-er(-est)
quasiequiv-a-lence  qua-si-equiv-a-lence

or quasi-

quasi-hy-ponor-mal  qua-si-hy-po-nor-mal
quasir-ad-i-cal  qua-si-rad-i-cal
quasiresid-ual  qua-si-resid-ual

qua-sis-mooth  qua-si-smooth
qua-sis-ta-tion-ary  qua-si-sta-tion-ary
qu-a-sito-pos  qua-si-topos
qu-a-si-tri-an-gu-lar  qua-si-tri-an-gu-lar

quintessence  quin-tes-sence
quintessen-tial  quin-tes-sen-tial

rab-bi-try  rab-bit-ry
ra-dio-g-ra-phy  ra-di-og-ra-phy
raf-f-ish(ly)  raff-ish(-ly)

ramshackle  ram-shackle
ravenous  rav-en-ous

re-ar-range-ment  re-arrange-ment

re-ciproc-i-ties  rec-i-proc-i-ties
reci-procity  rec-i-proc-i-ty

rect-an-gle  rec-tan-gle

+ re-ducible  re-duc-ible

ree-cho  re-echo

reprint(s)  re-print(s)
restorable  re-stor-able
retrofit(ted)  retro-fit(-ted)
+ reusable  re-us-able
+ reuse  re-use
rhinoceros  rhi-noc-er-os
righ-teous(ness)  right-eous(-ness)
ringleader  ring-leader
robot  ro-bot
robotics  ro-bot-ics
roundtable  round-table
salesclerk  sales-clerk

salescle-rks  sales-clerks

saleswoman(en)  sales-woman(en)
salmonella  sal-mo-nel-la
+ salta-tion  sal-ta-tion
sarsaparilla  sar-sa-par-il-la
sauerkraut  sauer-kraut

sca-to-log-i-cal  scat-o-log-i-cal

schedul-ing  sched-ul-ing
schizophrenic  schiz-o-phrenic
schnauzer  schnau-zer
schoolchild(ren)  school-child(-ren)
schoolteacher  school-teacher

scy-thing  scyth-ing

sec-re-tariat  sec-re-tar-iat
semaphore  sem-a-phore
semester  se-mes-ter
semidef-i-nite  semi-def-i-nite
+ semidi-rect  semi-di-rect

semi-ho-mo-th-etic  semi-ho-mo-thet-ic
semir-ing  semi-ring

+ semisim-ple  semi-sim-ple
semiskilled  semi-skilled
seroepi-demi-o-log-i-cal sero-epi-de-mi-o-log-i-cal

ser-vomech-a-nism  ser-vo-mech-anism

+ sesquipedalian  ses-qui-pe-da-lian
setup(s)  set-up(s)
severely  se-vere-ly

sha-peable  shape-able

shoestring  shoe-string
sidestep  side-step
sideswipe  side-swipe
skyscraper  sky-scraper
smokestack  smoke-stack

snorke-l-ing  snor-kel-ing

solenoid  so-le-noid

so-lute(s)  solute(s)

sovereign  sov-er-eign
spaces  spa-ces
specious  spe-cious
spelunker  spe-lunk-er
spendthrift  spend-thrift

spheroid(al)  spher-oid(-al)

sph-inges  sphin-ges

spi-cily  spic-i-ly
spinors  spin-ors
spokeswoman(en)  spokes-woman(en)
sportscast(er)  sports-cast(-er)
sportively  spor-tive-ly
sportswear  sports-wear
sportswriter  sports-writer
sprightlier  spright-lier
squeamish  squea-mish

stan-dalone  stand-alone

startling(ly)  star-tling(-ly)
statis-tics  sta-tis-tics
stealthily  stealth-ily
steeplechase  steeple-chase
stochas-tic  sto-chas-tic
strangeness  strange-ness
+ straphanger  strap-hanger
stratagem  strat-a-gem
stretchier  stretch-i-er
stronghold  strong-hold
strongest  strong-est
stupi-der(est)  stu-pid-er(-est)

+ sub-d-if-fer-en-tial  sub-dif-fer-en-tial

summable  sum-ma-ble

su-perego  super-ego
su-pere-gos  super-egos

supremacist  su-prema-cist
surveil-lance  sur-veil-lance

swim-m-ingly  swim-ming-ly

symp-tomatic  symp-to-matic
syn-chromesh  syn-chro-mesh
syn-chronous  syn-chro-nous
syn-chrotron  syn-chro-tron
talkative  talk-a-tive
tapestry(ies)  ta-pes-try(ies)
tarpaulin  tar-pau-lin

tele-g-ra-pher  te-leg-ra-pher

telekinetic  tele-ki-net-ic
teler-obotics  tele-ro-bot-ics
tem-porar-ily  tem-po-rar-ily
tenure  ten-ure
testbed  test-bed

tex-twidth  text-width
tha-la-mus  thal-a-mus
ther-moe-las-tic  ther-mo-elas-tic
times-tamp  time-stamp

toolkit  tool-kit

to-po-graph-i-cal  topo-graph-i-cal
to-ques  toques

traitorous  trai-tor-ous
transceiver  trans-ceiver
transgress  trans-gress
transver-sal(s)  trans-ver-sal(s)
transvestite  trans-ves-tite
traversable  tra-vers-a-ble
traver-sal(s)  tra-ver-sal(s)
treacheries  treach-eries
troubadour  trou-ba-dour
turkey  tur-key
turnaround  turn-around

ty-pal  typ-al

unattached  un-at-tached

unerringly  un-err-ing-ly
un-friendly(ier)  un-friend-ly(i-er)

va-guer  vaguer

vaudeville  vaude-ville

vi-cars  vic-ars
vil-lai-ness  vil-lain-ess

viviparous  vi-vip-a-rous
voiceprint  voice-print

vs-pace                    \vspace

wadding  wad-ding

wallflower  wall-flower
wastew-a-ter  waste-water
waveg-uide  wave-guide

wavelets  wave-lets

we-b-like  web-like

weeknight  week-night
wheelchair  wheel-chair
whichever  which-ever
whitesided  white-sided

whites-pace  white-space

widespread  wide-spread
wingspread  wing-spread
witchcraft  witch-craft
wordspac-ing  word-spac-ing
workhorse  work-horse
wraparound  wrap-around
wretched(ly)  wretch-ed(-ly)
yesteryear  yes-ter-year

Names and non-English words used in English text

al-ge-brais-che  al-ge-brai-sche

Al-legheny  Al-le-ghe-ny
Arkansas  Ar-kan-sas
Aus-tralasian  Aus-tral-asian
au-toma-tisierter  auto-mati-sier-ter

Be-di-enung  Be-die-nung
+ Be-mbo  Bembo

bib-li-ographis-che  bib-li-o-gra-phi-sche
Boston  Bos-ton

Brow-n-ian  Brown-ian

Brunswick  Bruns-wick
Bu-dapest  Bu-da-pest
Caribbean  Car-ib-bean
Charleston  Charles-ton
Char-lottesville  Char-lottes-ville
Columbia  Co-lum-bia
Czechoslo-vakia  Czecho-slo-va-kia

Di-jk-stra  Dijk-stra

dy-namis-che  dy-na-mi-sche

En-glish  Eng-lish
Eu-le-rian  Euler-ian

Evanston  Evan-ston
Febru-ary  Feb-ru-ary
Festschrift  Fest-schrift
Florida  Flor-i-da
Forschungsin-sti-tut  For-schungs-in-sti-tut

funk-t-sional  funk-tsional
Gaus-sian  Gauss-ian
Greif-swald  Greifs-wald

Grothendieck  Grothen-dieck
Grundlehren  Grund-leh-ren
Haifa  Hai-fa

                                                          Hamil-to-nian  Hamil-ton-ian

Helsinki  Hel-sinki

                                                          Her-mi-tian  Her-mit-ian
                                                          Hi-bbs  Hibbs

Hokkaido  Hok-kai-do
Jan-uary  Jan-u-ary
Japanese  Ja-pa-nese
Kadomt-sev  Kad-om-tsev
Kansas  Kan-sas

                                                          Karl-sruhe  Karls-ruhe

Ko-rteweg  Kor-te-weg
Lan-caster  Lan-cas-ter

                                                          Leg-en-dre  Le-gendre
                                                          Le-ices-ter  Leices-ter
                                                          Lip-s-chitz(ian)  Lip-schitz(-ian)

Louisiana  Lou-i-si-ana
Manch-ester  Man-ches-ter
Marko-vian  Mar-kov-ian
Mas-sachusetts  Mass-a-chu-setts
Min-neapo-lis  Min-ne-ap-o-lis
Min-nesota  Min-ne-sota
Moscow  Mos-cow
Nachrichten  Nach-richten
Nashville  Nash-ville
Ni-jmegen  Nij-me-gen
Noethe-rian  Noe-ther-ian

                                                          No-ord-wi-jk-er-hout  Noord-wijker-hout

Novem-ber  No-vem-ber
Palermo  Pa-ler-mo
Philadel-phia  Phil-a-del-phia
Poincare  Poin-care

                                                          Po-ten-tial-gle-ichung  Po-ten-tial-glei-chung

rathskeller  raths-kel-ler

                                                          Rie-man-nian  Rie-mann-ian
                                                          Ry-d-berg  Ryd-berg
                                                          schot-tis-che  schot-tische

Schrodinger  Schro-ding-er
Schwabacher  Schwa-ba-cher
Schwarzschild  Schwarz-schild
Septem-ber  Sep-tem-ber

                                                          Stokess-che  Stokes-sche

Stuttgart  Stutt-gart
Susque-hanna  Sus-que-han-na
Taube-rian  Tau-ber-ian

                                                          tech-nis-che  tech-ni-sche

Ten-nessee  Ten-nes-see
ty-pographique  ty-po-graphique
Ukrainian  Ukrain-ian

                                                          ve-r-all-ge-mein-erte  ver-all-ge-mein-erte

Vere-ini-gung  Ver-ein-i-gung
Verteilun-gen  Ver-tei-lun-gen

                                                          Wahrschein-lichkeit-s-the-o-rie

Wahr-schein-lich-keits-theo-rie

Werthe-rian  Wer-ther-ian
Winch-ester  Win-ches-ter
+ Wirtschaft  Wirt-schaft
Yingy-ong Shuxue Jisuan  Ying-yong Shu-xue Ji-suan
Zealand  Zea-land
Zeitschrift  Zeit-schrift

⋄ Barbara Beeton

                                                                     bnb@Math.AMS.com